Angela is training for a marathon and completes her long mileage runs for training on the weekend. Over the last 3 weekends she ran 15 miles in 2 hours; 18 miles in 2 hours, 33 minutes; and 22 miles in 3 hours, 7 minutes. Determine if her weekend training runs showcase a proportional relationship.

Found 2 solutions by Theo, greenestamps:
Answer by Theo(13342)

the ratios are:
miles / minutes = 18 / 153 for the 18 mile run and 22 / 187 for the 22 mile run.
if it was a proportional relationship, then the ratios would be the same.
153/18 = 8.5
22/187 = 8.5
the ratios are the same, indicating a proportional relationship.
18/153 = .1176470588
22/187 = .1176470588
the ratios are the same.
another way is to use your calculator to get 18/153 - 22/187.
if the result is 0, they are the same ratio.
you could also use the old method of cross multiplication.
if the cross multipliers are equal, then the ratios are the same.
you should be:
18*187 = 22*153 which becomes 3366 = 3366.
they're the same, so the ratios are equivalent.




Answer by greenestamps(13200)



The data for the last two weekends show a proportional relationship. In miles per minute,
2nd weekend: 18/153 = 2/17
3rd weekend: 22/187 = 2/17

But the data for the first weekend shows a different rate: 15/120 = 1/8

ANSWER: NO, the training runs do not show an exactly proportional relationship.